0 0
Add a comment Improve this question Transcribed image text
Answer #1

Power output P =5MW, =>

heat rate Q = 5*10^6/0.35 =14.2857*10^6 W

b) Mass Energy Density =50 MJ/kg

=> amount methane needed = m(kg/s) * 50*10^6 J/kg =14.2857*10^6 J/s

=> m (kg/s) =14.2857*10^6/(50*10^6 ) =0.2857 kg/s = 0.2857 kg/(1s*1hr/3600s) =1028.5704 kg/hr